What and Define Equivalent formula or laws of algebra of propositions

1
Expert's answer
2021-11-29T16:24:48-0500

If

"\begin{aligned}
&p \leftrightarrow q \equiv(p \rightarrow q) \wedge(q \rightarrow p) \\
&p \leftrightarrow q \equiv(p \wedge q) \vee(\neg p \vee \neg p)
\end{aligned}"

Then

"\begin{gathered}
(\neg p \vee q) \wedge(p \vee \neg q)(\text { Commutation }) \\
(\neg p \vee q) \wedge(\neg q \vee p) \text { (Implication) } \\
(p \rightarrow q) \wedge(q \rightarrow p) \text { (Equivalence) } \\
\qquad \begin{aligned}
p & \leftrightarrow q \equiv(\text { Equivalence }) \\
(p \wedge q) & \vee(\neg p \wedge \neg q)(\text { DeMorgan }) \\
(p \wedge q) & \vee \neg(p \vee q)(\text { Commutation }) \\
\neg(p \wedge q) & \vee(p \wedge q)(\text { Implication }) \\
&(p \wedge q) \rightarrow(p \wedge q)
\end{aligned}
\end{gathered}"